# An Analytical and AI-discovered Stable, Accurate, and Generalizable Subgrid-scale Closure for Geophysical Turbulence

### v2

@@ Abstract
The “structural modeling” approach [2], e.g., based on truncated Taylor-series expansion of the SGS flux, yields closures such as the nonlinear gradient model (NGM2), which satisfies (1) with above 0.9 pattern correlation with filtered DNS [15, 16, 17]. However, these closures fail (2)-(3), e.g., leading to unstable LES [18, 19, 20, 21]. The “functional modeling” approach [22, 23, 24, 25] leads to eddy-viscosity closures such as Smagorinsky [26] and Leith [25], which produce stable LES and some aspects of (2) though they are often overly dissipative and miss backscattering (the transfer of energy from the SGS processes to the resolved scales), an important process in atmospheric and oceanic flows [27, 28, 29, 30, 31, 32, 4, 33, 34, 35, 36]. Excessive dissipation also degrades the representation of extreme events [37, 38]. Furthermore, these closures fail (1), e.g., lower than 0.5 pattern correlation [39, 37, 40].
@@
Recently, AI approaches for developing data-driven closures have received significant attention. Most efforts have focused on supervised learning with filtered DNS data, which is an approach akin to structural modeling [7, 41, 11, 42]. While some studies have demonstrated that these data-driven closures can satisfy conditions (1)–(3), they suffer from three major limitations: (i) the need for large amounts of DNS snapshots for training, (ii) lack of interpretability, and (iii) poor generalization to out-of-distribution regimes, e.g., extrapolation to much higher Reynolds numbers [43, 44, 37, 45, 46]. Other promising AI frameworks have also shown success, such as those that directly learn the coarse-grained dynamics with neural operators, or self-supervised approaches, akin to functional modeling, which use only DNS statistics [47, 48, 49, 50]. However, they can still suffer from shortcomings (ii) and (iii) .
@@
A number of studies have pursued an alternative approach: “equation discovery”, a class of AI techniques that aim to find a closed-form mathematical representation of the data [51, 52, 53, 54, 55, 56, 57, 58]. The major appeals of this approach are data efficiency, interpretability, and generalizability, thus addressing the shortcomings of deep learning. In their pioneering work, Zanna and Bolton [20] used Bayesian sparse regression to find a closed-form equation for SGS momentum stress in geophysical turbulence. They robustly discovered a closure that resembled NGM2 but found that while this closure accurately captured the detailed structure of the SGS fluxes, the LES was unstable (unless the predicted SGS flux was attenuated). Using the same methodology, Jakhar et al. [21] discovered closures of the same structure for SGS momentum and heat fluxes across various setups of 2D turbulence and Rayleigh-Bénard convection and confirmed that they all exactly match the outcome of second-order truncated Taylor-series expansion; i.e., the discovered closure is NGM2. They also demonstrated that the LES instability results from NGM2’s inability to capture any inter-scale energy transfer (neither diffusion nor backscattering). However, it remained unclear how a better equation discovery approach can be devised, though the need for a physics-informed loss function and a better library (e.g., through genetic programming [51, 33]) was speculated.
@@
The dimensionless continuity and momentum equations for LES of 2D turbulence in spatial dimensions ($x,y$) are given by:
@@
$\frac{\partial\overline{u}_{i}}{\partial t}+\frac{\partial\overline{u}_{i}\overline{u}_{j}}{\partial x_{j}}$ $=$ $-\frac{\partial\overline{p}}{\partial x_{i}}+\frac{1}{Re}\frac{\partial^{2}\overline{u}_{i}}{\partial x_{j}x_{j}}-\frac{\partial\tau_{ij}}{\partial x_{j}}+\overline{\mathcal{Q}}_{i}.$ (2)
Here, $\overline{(.)}$ represents a low-pass filtering operation, $u_{i}$ denotes velocity, $p$ is pressure, $Re$ is the Reynolds number, $\overline{\mathcal{Q}}_{i}$ represents a time-invariant external forcing, Rayleigh drag, and the Coriolis force. The SGS stress tensor, $\tau_{ij}=\overline{u_{i}u}_{j}-\overline{u}_{i}\overline{u}_{j}$, requires a closure model that expresses $\tau_{11},\tau_{12}=\tau_{21},$ and $\tau_{22},$ as a function of the resolved flow variables ($\overline{u}_{i},\overline{p}$).
We investigate four cases of 2D turbulence (see Figure 4), generating a diverse flow dynamics that vary in dominant length scales and energy ($E=\frac{1}{2}u_{i}u_{i}$) and enstrophy ($Z=\frac{1}{2}\omega^{2}$) cascade regimes ($\omega$ is the vorticity). Here, we regard direct numerical simulation (DNS) data as the ground truth and utilize filtered DNS (FDNS) data to learn closure for $\tau_{ij}$. The FDNS data are obtained by first filtering the DNS data using a Gaussian filter and then coarse-graining the results to the $4$ to $64\times$ coarser LES grid following Jakhar et al. [21].
The equation discovery method used in this study is a sparsity-promoting Bayesian linear regression approach [54, 20, 59] based on the relevance vector machine (RVM) [60], employed to derive closed-form closures for each element of $\tau_{ij}$ using FDNS data. This method operates on a library of basis functions, $\bm{\Phi}$, comprising linear or nonlinear combinations of relevant variables, such as velocity or its derivatives:
@@
where $A,B=\overline{u}_{1}$ and $\overline{u}_{2}$, and $C=\overline{\omega}$ or an element of $\overline{S}_{ij}=\frac{1}{2}\left(\frac{\partial\overline{u}_{i}}{\partial x_{j}}+\frac{\partial\overline{u}_{j}}{\partial y_{i}}\right)$, the strain rate. Material derivative $\frac{D}{Dt}=\frac{\partial}{\partial t}+u_{j}\frac{\partial}{\partial x_{j}}$ is included to account for non-Markovian (memory) effects, motivated by Mori-Zwanzig formalism [61, 62].
@@
For each of the four cases in Fig. 4, we separately discover closures for three elements of the SGS stress tensor. We also examine several LES grid sizes, and the effect of varying $\alpha$, which as mentioned earlier, is a key hyper‐parameter in equation discovery. We use $n$ = 100 FDNS samples from a training set and 20 FDNS samples from an independent testing set.
As an illustrative example, Fig. 1(a) displays the mean CC for all elements of $\tau_{ij}$ as $\alpha$ increases. Note that this structural modeling approach is what all past equation-discovery studies followed [20, 21]. For small $\alpha$ (<1<1), no closure is discovered (CC = 0, zero terms). As $\alpha$ increases, the number of discovered terms increases, and the CC value rises and eventually plateaus, forming an “L-Curve.” The elbow of this curve is commonly used to identify the $\alpha$ that balances accuracy and model size in equation discovery [54, 20, 59]. In Fig. 1(a), the CC-$\alpha$ relationship for each element of $\tau_{ij}$ converges to approximately 1 and robustly discovers the same model at the elbow. Past studies [20, 21] found that the elbow of this curve (red circles) corresponds to the analytically derivable NGM2 [16, 15]. However, as these studies and earlier work found, LES with NGM2 leads to unstable a posteriori (online) simulations. Jakhar et al. [21] suggested that this instability is due to the inability of NGM2 in capturing any inter-scale energy transfers; i.e., $P_{\tau}^{\text{NGM2}}(x,y)=0$ (NGM2 captures neither diffusion nor backscattering).
@@
It is noteworthy that the L-curves in Fig. 1 are consistently observed for all cases with any other $N_{\text{LES}}$. Note that increasing $\alpha$ leads to another closure with more terms but negligible improvements in CC; we have found this closure to be NGM6; see Eq. (4) (a generalized Taylor-series expansion of $\tau_{ij}$ is provided in the supplementary information). The equation discovery with the expansive 930-term library confirms that Taylor-series expansion offers the best representation of both $\tau_{ij}$ and the inter-scale energy transfer. Note that the memory terms never emerged during the discovery.
@@
Figure 2: Comparison of the a priori (offline) performance of different closures. The table shows the ratio of domain-averaged $P_{\tau}$ and $P_{Z}$ to that of FDNS. The bar plots present the CC for $\tau_{12}$, $P_{\tau}$, and $P_{Z}$ of each closure and FDNS. All values are the mean and standard deviations calculated for 4 setups together (Cases 1-2 at $N_{\text{LES}}=128$ and Cases 3-4 at $N_{\text{LES}}=512$); Table S1 shows the values for each case.
We first compare the a priori (offline) performance of NGMs (NGM2-6) with common baseline closures, the eddy-viscosity Smagorinsky (Smag) [26], its dynamic version (DSmag) [23], and the dynamic Leith (DLeith) [25] models (see Appendix B). Figure 2 shows that NGM4 and NGM6 perfectly capture the structure of the FDNS SGS stress and inter-scale enstrophy transfers, and very well capture the inter-scale energy transfers (both diffusion and backscattering), significantly outperforming NGM2 and the physics-based closures.
An important feature of NGM4 and NGM6 is that they capture the energy backscatter that is seen in the FDNS. While the overall representations of inter-scale energy transfer of NGM4 and NGM6 are not perfect (around $62\%$ and $78\%$ of FDNS’, respectively), they are substantially better than the energy transfers of NGM2 (=0=0) and the eddy-viscosity closures, which are purely and excessively diffusive (as indicated by the negative CC values). The advantages of NGM4 and 6, especially over the eddy-viscosity closures, can also be seen in the inter-scale energy and enstrophy transfers’ spectra (Fig. S1). These a priori tests show that NGM4 and NGM6 are fairly comparable, though NGM6 has a better $P_{\tau}$ (see Table S1).
Figure 3: Comparison of the a posteriori performance of different closures for Cases 3-4 at $N_{\text{LES}}=512$. (a)-(b) The PDF of vorticity normalized by its standard deviation, $\overline{\omega}/\sigma_{\omega}$. The upper-right insets of (a)-(b) compares the accuracy of short-term forecasting of each LES against FDNS. The $x$-axis represents the eddy-turnover time, $t_{\eta}=1/\sqrt{\langle\overline{\omega}^{2}\rangle}$. (c)-(d) Comparison of the enstrophy spectra, $\hat{Z}\left(k\right)$, of LES with different closures. A similar figure for (Cases 1-2 at $N_{\text{LES}}=128$ ) is shown in the supporting information.
@@
As shown in Figs. 3 and S3, when stable, LES with NGM4 and NGM6 capture both the bulk and tails of the probability distribution function (PDF) of vorticity (the tails of this PDF represent “extreme weather” in this prototype of atmospheric circulation). LES with eddy-viscosity closures fail to capture the extremes, especially for Cases 3 and 4, which have high-wavenumber forcings. Note that NGM2 for any case and any LES resolution leads to instabilities.
@@
In addition to better capturing extreme events, LES with NGM4 and NGM6 outperforms eddy-viscosity models in short-term forecasting of vorticity (see insets of panels (a)-(b) in Figs. 3 and S3). LES with NGM4 and NGM6 also outperform LES with other closures in capturing the domain-averaged kinetic energy and enstrophy (Table S2). Figures 3 and S3 (c)-(d) show the enstrophy spectra of FDNS and LES. The insets magnifying the tails of the spectra do not clearly show the superiority of LES with any specific closure. However, the insets magnifying the spectra at lower wavenumbers reveal that LES with eddy-viscosity closures underestimate the enstrophy at large scales, likely due to the excessive diffusion. The same observation can be made about the energy spectra (Fig. S4). Note that in a posteriori tests, LES with NGM4 and NGM6 show fairly similar performances, suggesting that in practice, using NGM4 might be enough.
@@
In conclusion, a physics-informed data-driven discovery identified an analytically derivable, yet previously ignored, representation of SGS closure that satisfies both structural and functional modeling requirements. This closure (NGM4) is “interpretable” as it is derivable from the Taylor-series expansion and it represents the Leonard and cross stresses (see supporting information). This closure is “generalizable” since its coefficients are entirely determined by the filter width, ($\Delta$) and does not have any component that depends on $Re$ or any other parameters. NGM4 is accurate: it is the first closure to simultaneously excel in both a priori and a posteriori tests. LES with NGM4 is stable only when the numerical resolution is high enough to capture $80\%$ of DNS enstrophy; however, this is consistent with general limitations of LES [1].
@@
Although this closure could have been derived analytically decades ago, AI played a critical role in its discovery here. First, the analytical expansion of the gradient model to higher-order terms had not been previously pursued. Among the main reasons are that the source of the shortcoming of NGM2 was unclear and gradient closures were generally ignored in favor of stable eddy-viscosity ones. The sparse equation discovery showed that among 930 terms, NGM4 and NGM6 best represent both the stress tensor and inter-scale energy transfer, and that commonly discussed phenomena such as eddy-viscosity and temporal memory are not needed. It was only after the AI-based discovery that the connection with higher-order Taylor-series expansion became clear.
@@
A major difference between NGM4 and NGM6 is that the latter represents the Reynolds stress too (see supporting information). However, while a priori tests show that NGM6 better represents the inter-scale energy transfer, a posteriori tests do not demonstrate any advantage over NGM4 in terms of stability or accuracy. This is encouraging as implementing models with high-order derivatives in numerical solvers can be challenging; these findings suggest that NGM4 might suffice in practice, e.g., in ocean modeling.
We note that NGM4 and NGM6 violate the Boussinesq hypothesis as they depend on second-order velocity gradients, highlighting the need to go beyond classical turbulence tensors to represent backscatter and anisotropic effects [64].
@@
While here we only focused on the SGS stress tensor in 2D turbulence, these findings are relevant to multi-scale modeling in other dynamical systems too, as coarse-graining any quadratic nonlinearity can lead to an NGM-like representation, e.g., as already shown for heat flux [21]. This work also presents the importance of combining physical and mathematical insight to better harness the power of AI methods in accelerating scientific discovery.

### v3

@@ Abstract
The “structural modeling” approach [2], e.g., based on truncated Taylor-series expansion of the SGS flux, yields closures such as the nonlinear gradient model (NGM2), which satisfies (1) with above 0.9 pattern correlation with filtered DNS (FDNS) [15, 16, 17]. However, these closures fail (2)-(3), e.g., producing no inter-scale energy transfer in 2D turbulence and leading to unstable LES [18, 19, 20, 21]. The “functional modeling” approach [22, 23, 24, 25] leads to eddy-viscosity closures such as Smagorinsky [26] and Leith [25], which produce stable LES and some aspects of (2) though they are often overly dissipative and miss backscattering (the transfer of energy from the SGS processes to the resolved scales), an important process in atmospheric and oceanic flows [27, 28, 29, 30, 31, 32, 4, 33, 34, 35, 36]. Excessive dissipation also degrades the representation of extreme events [37, 38]. Furthermore, these closures fail (1), e.g., lower than 0.5 pattern correlation [39, 37, 40].
@@
Recently, AI approaches for developing data-driven closures have received significant attention. Most efforts have focused on supervised learning with FDNS data, which is an approach akin to structural modeling [7, 41, 11, 42]. While some studies have demonstrated that these data-driven closures can satisfy conditions (1)–(3), they suffer from three major limitations: (i) the need for large amounts of DNS snapshots for training, (ii) lack of interpretability, and (iii) poor generalization to out-of-distribution regimes, e.g., extrapolation to much higher Reynolds numbers [43, 44, 37, 45, 46]. Other AI frameworks have also shown success, including self-supervised approaches, akin to functional modeling, which use only DNS statistics, and those that directly learn the coarse-grained dynamics [47, 48, 49, 50]. However, these approaches can still suffer from shortcomings (ii) and (iii).
@@
A number of studies have pursued an alternative approach: “equation discovery”, a class of AI techniques that aim to find a closed-form mathematical representation of the data [51, 52, 53, 54, 55, 56, 57, 58]. The major appeals of this approach are data efficiency, interpretability, and generalizability, thus addressing the shortcomings of deep learning. In their pioneering work, Zanna and Bolton [20] used Bayesian sparse regression to find a closed-form equation for SGS momentum stress in geophysical turbulence. They robustly discovered a closure that resembled NGM2 but found that while this closure accurately captured the detailed structure of the SGS fluxes, the LES was unstable (unless the predicted SGS flux was attenuated). Using the same methodology, [21] discovered closures of the same structure for SGS momentum and heat fluxes across various setups of 2D turbulence and Rayleigh-Bénard convection and confirmed that they all exactly match the outcome of second-order truncated Taylor-series expansion; i.e., the discovered closure is NGM2. They also demonstrated that the LES instability results from NGM2’s inability to capture any inter-scale energy transfer (neither diffusion nor backscattering). However, it remained unclear how a better equation discovery approach can be devised, though the need for a physics-informed loss function and a better library (e.g., through genetic programming [51, 33]) was speculated.
@@
The dimensionless continuity and momentum equations for LES of 2D turbulence are given by:
@@
$\frac{\partial\overline{u}_{i}}{\partial t}+\frac{\partial\overline{u}_{i}\overline{u}_{j}}{\partial x_{j}}$ $=$ $-\frac{\partial\overline{p}}{\partial x_{i}}+\frac{1}{Re}\frac{\partial^{2}\overline{u}_{i}}{\partial x_{j}\partial x_{j}}-\frac{\partial\tau_{ij}}{\partial x_{j}}+\overline{\mathcal{Q}}_{i}.$ (2)
Here, $\overline{(.)}$ represents a low-pass filtering operation, $u_{i}$ denotes velocity, $p$ is pressure, $Re$ is the Reynolds number, and $\overline{\mathcal{Q}}_{i}$ represents a time-invariant external forcing, Rayleigh drag, and the Coriolis force. The SGS stress tensor, $\tau_{ij}=\overline{u_{i}u}_{j}-\overline{u}_{i}\overline{u}_{j}$, requires a closure model that expresses $\tau_{11},\tau_{12}=\tau_{21},$ and $\tau_{22},$ as a function of the resolved flow variables ($\overline{u}_{i},\overline{p}$).
We investigate four cases of 2D turbulence, generating a diverse flow dynamics that vary in dominant length scales and energy ($E=\frac{1}{2}u_{i}u_{i}$) and enstrophy ($Z=\frac{1}{2}\omega^{2}$) cascade regimes ($\omega$ is the vorticity); see Appendix A for more information about these cases and their properties. Here, we regard direct numerical simulation (DNS) data as the ground truth (details of the numerical solvers are in Appendix C). We utilize FDNS data to learn closure for $\tau_{ij}$. The FDNS data are obtained by first filtering the DNS data using a Gaussian filter and then coarse-graining the results to the $4$ to $64\times$ coarser LES grid following Jakhar et al. [21]. The equation discovery method used in this study is a sparsity-promoting Bayesian linear regression approach [54, 20, 59] based on the relevance vector machine (RVM) [60], employed to derive closed-form closures for each element of $\tau_{ij}$ using FDNS data. This method operates on a library of basis functions, $\bm{\Phi}$, comprising linear or nonlinear combinations of relevant variables, such as velocity or its derivatives:
@@
where $A,B=\overline{u}_{1}$ and $\overline{u}_{2}$, $(x,y)=(x_{1},x_{2})$, and $C=\overline{\omega}$ or an element of $\overline{S}_{ij}=\frac{1}{2}\left(\frac{\partial\overline{u}_{i}}{\partial x_{j}}+\frac{\partial\overline{u}_{j}}{\partial x_{i}}\right)$, the strain rate. Material derivative $\frac{D}{Dt}=\frac{\partial}{\partial t}+u_{j}\frac{\partial}{\partial x_{j}}$ is included to account for non-Markovian (memory) effects, motivated by Mori-Zwanzig formalism [61, 62].
@@
For each of the four cases in Fig. 4, we separately discover closures for three elements of the SGS stress tensor. We also examine several LES grid sizes, and the effect of varying $\alpha$, which as mentioned earlier, is a key hyper‐parameter in equation discovery. We use $n$ = 100 FDNS samples from a training set and 20 FDNS samples from an independent testing set. We evaluate the a priori performance using the commonly employed average correlation coefficient (CC), Eq. (5), on the test samples [2, 71, 72, 21].
As an illustrative example, Fig. 1(a) displays the mean CC for all elements of $\tau_{ij}$ as $\alpha$ increases. Note that this structural modeling approach is what all past equation-discovery studies followed [20, 21]. For small $\alpha$ (<1<1), no closure is discovered (CC = 0, zero terms). As $\alpha$ increases, the number of discovered terms increases, and the CC value rises and eventually plateaus, forming an “L-Curve.” The elbow of this curve is commonly used to identify the $\alpha$ that balances accuracy and model size in equation discovery [54, 20, 59]. In Fig. 1(a), the CC-$\alpha$ relationship for each element of $\tau_{ij}$ converges to approximately 1 and robustly discovers the same model at the elbow. Past studies [20, 21] found that the elbow of this curve (red circles) corresponds to the analytically derivable NGM2 [16, 15]. However, as these studies and earlier work found, LES with NGM2 leads to unstable a posteriori (online) simulations. Jakhar et al. [21] suggested that this instability is due to the inability of NGM2 in capturing any inter-scale energy transfers; i.e., $P_{\tau}^{\text{NGM2}}(x_{1},x_{2})=0$ (NGM2 captures neither diffusion nor backscattering).
@@
It is noteworthy that the L-curves in Fig. 1 are consistently observed for all cases with any other $N_{\text{LES}}$. Note that increasing $\alpha$ leads to another closure with more terms but negligible improvements in CC; we have found this closure to be NGM6; see Eq. (4) (a generalized Taylor-series expansion of $\tau_{ij}$ is provided in the supplementary information [73]).
@@
The equation discovery with the expansive 930-term library confirms that Taylor-series expansion offers the best representation of both $\tau_{ij}$ and the inter-scale energy transfer. Note that the memory terms never emerged during the discovery.
Figure 2: Comparison of the a priori (offline) performance of different closures. The table shows the ratio of domain-averaged $P_{\tau}$ and $P_{Z}$ to that of FDNS. The bar plots present the CC for $\tau_{12}$, $P_{\tau}$, and $P_{Z}$ of each closure and FDNS. All values are the mean and standard deviations calculated for the 4 setups together (Cases 1-2 at $N_{\text{LES}}=128$ and Cases 3-4 at $N_{\text{LES}}=512$); Table S1 shows the values for each case.
We first compare the a priori (offline) performance of NGMs (NGM2-6) with common baseline closures, the eddy-viscosity Smagorinsky (Smag) [26], its dynamic version (DSmag) [23], and the dynamic Leith (DLeith) [25] models (see Appendix D). Figure 2 shows that NGM4 and NGM6 perfectly capture the structure of the FDNS SGS stress and inter-scale enstrophy transfers, and very well capture the inter-scale energy transfers (both diffusion and backscattering), significantly outperforming NGM2 and the physics-based closures.
An important feature of NGM4 and NGM6 is that they capture the energy backscatter that is seen in the FDNS. While the overall representations of inter-scale energy transfer of NGM4 and NGM6 are not perfect (around $62\%$ and $78\%$ of FDNS, respectively), they are substantially better than the energy transfers of NGM2 (=0=0) and the eddy-viscosity closures, which are purely and excessively diffusive (as indicated by the negative CC values). The advantages of NGM4 and 6, especially over the eddy-viscosity closures, can also be seen in the inter-scale energy and enstrophy transfers’ spectra (Fig. S1). These a priori tests show that NGM4 and NGM6 are fairly comparable, though NGM6 has a better $P_{\tau}$ (see Table S1).
Figure 3: Comparison of the a posteriori performance of different closures for Cases 3-4 at $N_{\text{LES}}=512$. (a)-(b) The PDF of vorticity normalized by its standard deviation, $\overline{\omega}/\sigma_{\omega}$. The upper-right insets of (a)-(b) compares the accuracy of short-term forecasting of each LES against FDNS. The $x$-axis represents the eddy-turnover time, $t_{\eta}=1/\sqrt{\langle\overline{\omega}^{2}\rangle}$. (c)-(d) Comparison of the enstrophy spectra, $\hat{Z}\left(k\right)$, of LES with different closures. A similar figure for (Cases 1-2 at $N_{\text{LES}}=128$ and Case 4 at $N_{\text{LES}}=1024$) is shown in the supplementary information.
@@
In the atmosphere, intense cyclonic and anticyclonic vortices, which have positive and negative vorticity and are associated with low- and high-pressure systems, respectively, can drive various kinds of weather extremes [75]. Therefore, in this prototype of atmospheric turbulent circulation, the tails of the probability distribution function (PDF) of vorticity represent low-probability, rare “extreme weather” events. As shown in Figs. 3 and S3-S4, when stable, LES with NGM4 and NGM6 capture both the bulk and tails of the PDF of vorticity. LES with eddy-viscosity closures fail to capture the extremes, especially for Cases 3 and 4, which have high-wavenumber forcings. Note that using NGM2 for any case and any LES resolution leads to instabilities.
In addition to better capturing extreme events, LES with NGM4 and NGM6 outperforms eddy-viscosity models in short-term forecasting of vorticity (see insets of panels (a)-(b) in Figs. 3 and S3-S4).
@@
LES with NGM4 and NGM6 also outperform LES with other closures in capturing the domain-averaged kinetic energy and enstrophy (Table S2). Figures 3, S3 (c)-(d), and S4 (b) show the enstrophy spectra of FDNS and LES. The insets magnifying the tails of the spectra do not clearly show the superiority of LES with any specific closure. However, the insets magnifying the spectra at lower wavenumbers reveal that LES with eddy-viscosity closures underestimate the enstrophy at large scales, likely due to the excessive diffusion. The same observation can be made about the energy spectra (Fig. S5). Note that in a posteriori tests, LES with NGM4 and NGM6 show fairly similar performances, suggesting that in practice, using NGM4 might be enough.
@@
In conclusion, a physics-informed data-driven discovery identified an analytically derivable, yet previously ignored, representation of SGS closure that satisfies both structural and functional modeling requirements. This closure (NGM4) is “interpretable” as it is derivable from the Taylor-series expansion and it represents the Leonard and cross stresses (see supplementary information [73]). This closure is “generalizable” since its coefficients are entirely determined by the filter width ($\Delta$) and does not have any component that depends on $Re$ or any other parameters. NGM4 is accurate: it is the first closure to simultaneously excel in both a priori and a posteriori tests. LES with NGM4 is stable only when the numerical resolution is high enough to capture $80\%$ of DNS enstrophy; however, this is consistent with general limitations of (explicit) LES [1] (see below for more discussions).
@@
Although this closure could have been derived analytically decades ago, AI played a critical role in its discovery here. First, the analytical expansion of the gradient model to higher-order terms had not been previously pursued. Among the main reasons are that the source of the shortcoming of NGM2 was unclear and gradient closures were generally ignored in favor of stable eddy-viscosity ones. The sparse equation discovery showed that among 930 terms, NGM4 and NGM6 best represent both the stress tensor and inter-scale energy transfer. It was only after the AI-based discovery that the connection with higher-order Taylor-series expansion became clear.
@@
A major difference between NGM4 and NGM6 is that the latter represents the Reynolds stress too (see supplementary information [73]). However, while a priori tests show that NGM6 better represents the inter-scale energy transfer, a posteriori tests do not demonstrate any advantage over NGM4 in terms of stability or accuracy. This is encouraging as implementing models with high-order derivatives in numerical solvers can be challenging; these findings suggest that NGM4 might suffice in practice, e.g., in ocean modeling. We note that NGM4 and NGM6 violate the Boussinesq hypothesis as they depend on second-order velocity gradients, highlighting the need to go beyond classical turbulence tensors to represent backscatter and anisotropic effects [64].
The findings of this work can also be framed within the broader context of nonlocal closures [76, 77, 78, 79, 80]. The finite-order gradient closures (Eq. (4)) can be viewed as truncations of the nonlocal deconvolution operator [2]. Therefore, NGM4 can be interpreted as the minimal local truncation that successfully captures the leading-order nonlocal effects required by our reconstruction and inter-scale transfer criteria. This interpretation suggests that for cases requiring stronger scale-selectivity or wider spectral fidelity, Padé-like rational approximations [81] or explicitly nonlocal kernels can be a natural extension.
Furthermore, we find that NGM4, a deterministic, local closure, suffices for these cases of 2D turbulence with our chosen criteria. However, this still leaves the door open for stochastic (e.g., generative) extensions and closures with memory. In particular, applying equation-discovery methodologies to find parsimonious stochastic closures [82, 83] is therefore a promising direction for future work.
It is also important to note that in practice, many applications, such as weather and climate modeling, use “implicit” LES [84, 85]. In these highly under-resolved simulations, the numerical discretization acts as an implicit closure, effectively fusing the numerical method and the SGS modeling. Our work presents a testable, resolution-aware closure that, in the explicit LES regime, is compelling for its verified inter-scale energetics and a posteriori performance. In the under-resolved implicit LES regimes, this closure can serve as a target for diagnostics of inter-scale transfers and extremes, aid interpretability studies, and potentially be combined with existing discretization schemes when an explicit representation of backscatter is desired.
@@
Finally, it should be noted that while here we only focused on the SGS stress tensor in 2D turbulence, these findings are relevant to multi-scale modeling in other dynamical systems too, as coarse-graining any quadratic nonlinearity can lead to an NGM-like representation, e.g., as already shown for heat flux in Rayleigh-Bénard convection [21]. This work also presents the importance of combining physical and mathematical insight to better harness the power of AI methods in accelerating scientific discovery.